A straight chain of hydrocarbon has molecular formula C8H10 .The hybridization for the carbon atoms from one sp3 , sp2 ,sp2, sp3,sp2,sp2,sp and sp. The structural formula of hydrocarbon would be ?

Grade:11

1 Answers

Pooja
5 years ago
Dear Student,
Please calculate the degree of unsaturation in the given formula of the compound
which will come out to be 4.

CH3-CH=CH-CH2-CH=CH-C=-CH
there is a triple bond at the end.